OpenStudy (anonymous):

2/3x+40=x/2-x/6-2x/9...solve linear equation

OpenStudy (anonymous):

just want to clarify, its 2/(3x) or (2/3) x

OpenStudy (anonymous):

(2/3) x

OpenStudy (mertsj):

Add the three fractions on the right : x(1/2-1/6-2/9)=1/9x

OpenStudy (mertsj):

Subtract 2/3 x from both sides: 40=-5/9(x) Multiply both sides by -9/5 x=-72